OpenStudy (shaeelynn):

4Square root x^20y^28

OpenStudy (ilcapitano):

I... I assume you're being asked to simplify this expression? If so, you should think of the square root as a one-half power (I could explain why that makes sense if you want.) and then simplify the exponents. If I read it like, it would go like this: 4 * sqrt(x^20*y^28) = 4 * (x^20 * y^28)^(1/2) = 4 * (x^20)^(1/2) * (y^28)^(1/2) = 4 * x^(20/2) * y^(28/2) = 4 * x^10 * y^14

OpenStudy (shaeelynn):

no I know how to solve it with a normal 4 in front of the square root but in this case its a tiny 4

OpenStudy (ilcapitano):

A "tiny" four? Is it in the cup on the root operator? If so then it's a "fourth root" instead of a "square root". Fourth roots are "one-quarter" powers.

OpenStudy (shaeelynn):

yeah its the 4th root sorry i forgot what it was called

OpenStudy (ilcapitano):

So it's more or less the same process: \[\sqrt[4]{x^{20}y^{28}} = (x^{20}y^{28})^{(1/4)}\] ...and then simplify from there.

OpenStudy (shaeelynn):

yeah im having trouble with the simplifying part

OpenStudy (agent0smith):

Use exponent rules \[\Large (x^a)^b=x^{a*b}\]

jhonyy9 (jhonyy9):

yes exactly just multiplie the 20 by 1/4 and the 28 by 1/4 ok. ?

OpenStudy (shaeelynn):

oh so i just multiply?

OpenStudy (agent0smith):

The exponents, yes.

OpenStudy (shaeelynn):

so it would be 20^5y^7?

OpenStudy (agent0smith):

Why would it be 20^5?

OpenStudy (shaeelynn):

because 20(1/4) is 5

OpenStudy (agent0smith):

\[\Large (x^{20}y^{28})^{1/4}=\]

OpenStudy (shaeelynn):

\[x ^{5}y ^{7}\]

OpenStudy (agent0smith):

Much better.

OpenStudy (shaeelynn):

so is that the answer?

OpenStudy (agent0smith):

Well, do you see anything else you can do to it?

OpenStudy (shaeelynn):

no

OpenStudy (agent0smith):

Exactly.